The area of a triangle is 18 square meters, and the base of the triangle is 9 meters.

Find the altitude of the triangle.



1
Expert's answer
2021-01-28T05:09:50-0500
"Area=\dfrac{1}{2}\times base \times altitude"

"altitude=\dfrac{2\times Area}{base}"

"altitude=\dfrac{2\times 18m^2}{9m}=4m"

The altitude of the triangle is 4 meters.